mework: 9.1 re: O of 1 pt HYV Score: 3333%, 5 of 15 US FLAG
Answer #1

X = 534 , n = 1012

a)

p^= X/n = 534/1012 = 0.527667

= 0.5277

95% confidence interval (0.497,0.558)